OpenStudy (anonymous):

How do u solve this question using elimination m - n = -9 & 7m + 2n = 9? Please help me!!!

OpenStudy (texaschic101):

m - n = -9 -->(2) 7m + 2n = 9 ----------------- 2m - 2n = -18 (result of multiplying by 2) 7m + 2n = 9 -----------------add 9m = -9 m = -1 now sub -1 in for m in either of the original equations m - n = -9 -1 - n = -9 -n = -9 + 1 -n = -8 n = 8 check... 7m + 2n = 9 7(-1) + 2(8) = 9 -7 + 16 = 9 9 = 9 (correct) m = -1 and n = 8 any questions ?

OpenStudy (anonymous):

No and thanx :)

OpenStudy (texaschic101):

YW :)